Combinatorial Optimization - Presentation Review

Ehsan Moradi Pari

LCSEE, West Virginia University,

Morgantown, WV {ksmani@csee.wvu.edu}

April 2, 2013

1 Instructions

(a) Give a score between 1 and 5 for each question, and justify the same. Sum up the individual scores and provide the total on the top of the sheet.

(b) Please answer questions professionally and sincerely.

2 Review

1. Is the lecturer well-prepared?

2. Did the instructor answer your questions to your satisfaction?

3. Did the instructor maintain a good pace?

4. Did you understand the material communicated by the instructor?

5. Grade the the strengths of instructor.

6. Grade the weaknesses of instructor (5- no weakness. 0 - Very weak.)

7. Overall, how would you rate this lecture?

8. Overall, how would you rate this lecturer?

9. How much did you learn from this lecture?

10. How effective is the instructor?